OpenStudy (anonymous):

Help!!! Please!! In a hurry!! A stop sign is a regular octagon. This stop sign has a height of 30 inches and each side is 12.43 inches. The area of a 30-inch sign is ____ in2 to the nearest square inch.

OpenStudy (anonymous):

This is as simple as finding the area of a triangle. You have a right triangle from the top of the sign, with side b = 30/2 = 15 inches, and side A = 12.43/2 = 6.215 inches. Triangle area = 0.5ba = 0.5(15)(6.215) = 46.6125 inches. Now, you have exactly 2*8 of those triangles, two for each of 8 sides. The answer is 16*46.6125.

OpenStudy (anonymous):

THANK YOU VERY MUCH!!!

OpenStudy (anonymous):

yep no problem.

OpenStudy (anonymous):

I reallly appreciate it!!!

OpenStudy (anonymous):

Please give me a medal :)

OpenStudy (anonymous):

Okiedoki. When I figure out ho t do tht, I will indeed. I am still new this site, so I am still figuring things out.

OpenStudy (anonymous):

Just click the blue button at the top of my response. It's my first day here as well, good luck on your studies!~

OpenStudy (anonymous):

Oh, ok. Thank you again. And you too!